Show that ∞n=1 (x^n)/(n^n) converges for all x ∈ R


1
Expert's answer
2021-02-12T18:06:46-0500

Let us show that "\sum_{n=1}^{\infty} \frac{x^n}{n^n}" converges for all "x \in\mathbb R." Let us use Cauchy's root convergence test for series:

"L=\lim\limits_{n\to\infty}\sqrt[n]{|\frac{x^n}{n^n}|}=\lim\limits_{n\to\infty}\frac{|x|}{n}=|x|\cdot\lim\limits_{n\to\infty}\frac{1}{n}=|x|\cdot 0=0"


Since "L<1" for "x\in\mathbb R," we conclude that "\sum_{n=1}^{\infty} \frac{x^n}{n^n}" converges for all "x \in\mathbb R."